Circular No. 6083

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

KS 1730−312

K. Borozdin, N. Alexandrovich, V. Arefiev and R. Sunyaev, Space Research Institute, Moscow, on behalf of the MIR-KVANT-TTM team, communicate: “During observations of the Galactic Center field by the TTM coded-mask instrument onboard the MIR space station on Sept. 23, a flux of 70 ± 10 mCrab within the energy band 2–30 keV was detected from α = 17^{h}30^{m}.3, δ = −31^{o}11′ (equinox 1950.0, uncertainty ±2′). The source spectrum can be approximated by a power law with spectral index −1.9 ± 0.4.”

E. Churazov, M. Gilfanov, J. Ballet and E. Jourdain, on behalf of the Granat team (Space Research Institute, Moscow; Service d’Astrophysique, Centre d’Etudes de Saclay; and Centre d’Etude Spatiale des Rayonnements, Toulouse), report: “During observations of the Galactic Center region on Sept. 22-23 the SIGMA telescope onboard Granat Observatory detected a new hard x-ray source that is compatible with the above position. The 40–150-keV flux from the source was ∼ 130 mCrab. No significant emission from this source has been detected by SIGMA in previous observations, performed during Sept. 1–22 with 3σ upper limits of 40 mCrab (Sept. 20– 22) and 15 mCrab (Sept. 1–22). In subsequent observations on Sept. 24–25 and 25–26 the source flux was 116 and 93 mCrab, respectively. Gradual softening of the spectrum has been observed with the power-law photon index changing from 2.2 ± 0.2 on Sept. 22–23 to 2.9 ± 0.5 on Sept. 25–26.”

X-RAY NOVA 1993 IN OPHIUCHUS Churazov et al. add: “In the course of the same set of Galactic Center

observations SIGMA detected hard x-ray emission from the x-ray nova observed in Ophiuchus in 1993 (cf. IAUC 5874, 5876, et seq.; α = 17^{h}16^{m}.5, δ = −24^{o}58′, equinox 1950.0). The source was present in the 40–150-keV band during the whole interval 1994 Sept. 1–23. The time-averaged flux from the source is ∼ 110 mCrab. A power-law approximation to the source spectrum gives the photon index 2.3 ± 0.3.”

Borozdin and Alexandrovich report: “The TTM coded-mask instrument onboard MIR registered x-ray emission from the direction of the 1993 x-ray nova with flux 55 ± 15 mCrab (at 2–30 keV), averaged over observations on 1994 Aug. 31, Sept. 1 and 23. The photon index of the power-law approximation is 2.0 ± 0.7.”

1994 September 30 Brian G. Marsden